Circular No. 5840

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. Telephone 617-495-7244/7440/7444 (for emergency use only) TWX 710-320-6842 ASTROGRAM CAM EASYLINK 62794505 MARSDEN@CFA or GREEN@CFA (.SPAN, .BITNET or .HARVARD.EDU)

PERSEID METEORS With reference to the remark on IAUC 5839 concerning the possible

observation of a glow arising from sunlight scattered by meteoroidal dust, D. I. Steel, Anglo-Australian Observatory, draws attention to a discussion of this topic by Baggaley (1977, Obs. 97, 123), whose prediction in the case of the Perseids is very different from that given by Rao (1993, Sky Tel. 86, No. 2, 48). The point to watch must be the true radiant (and antiradiant), which has been computed by the undersigned to be at α = 33^{o}.5, δ = +84^{o}.7 (and its antipodal point); this differs by 27^{o} from the apparent (geocentric) radiant (i.e., that modified by the earth’s orbital velocity), which the current orbit of P/Swift-Tuttle indicates to be at α = 46^{o}.4, δ = +57^{o}.8 (equinox 2000.0). This computation is confirmed by Steel, and also by D. Kessler and H. Zook, NASA Johnson Space Center, who remark that, if there should be a strong concentration of material exactly such that it will collide with the earth, then the glow will be at the apparent radiant. They therefore suggest that, before the earth crosses the comet’s orbital plane, the arc between the apparent and true radiants should be monitored; as the orbital plane is crossed, any glow should converge on the true radiant—to jump to the true anti-radiant and then to the arc joining the true and apparent anti-radiants afterward. It seems that the glow points indicated by Rao considered the adjustment from apparent to true radiant with the wrong sign.

J. G. Hills and R. Hoffman, Los Alamos National Laboratory, discuss the possibility of observing larger individual pieces of cometary debris outside the earth’s atmosphere. For the purpose of searching for these bodies, one needs to concentrate on the apparent radiant and anti-radiant, as well as that half of the great circle joining them that crosses the equator near α = 136^{o}.4.

Further details regarding the ejection simulation by Williams and Wu (IAUC 5839) show that the basic assumption was that 1-mm particles were ejected from the comet in various directions from the sunward side at a speed of 50r^{−9/8} m/s, 99 percent of them when the comet was within r = 2 AU of the sun in 1862, forces due to planetary gravitational attractions and radiation pressure being considered. The suggestion that the Perseids might peak in 1994 was based on a cometary orbit that was too small. Use of the correct orbit shows the current activity starting around 1989 and suggests that the peak might not be until 1995 or even 1997.

1993 August 10 Brian G. Marsden